x+y-4=0 and 2x-y-2=0 awaiting the point (0,0) find the equation of a straight line
Answer by stanbon(75887)

x+y-4=0
2x-y-2=0
------
Add and solve for "x":
3x - 6 = 0
3x = 6
x = 2
------
Solve for "y":
x + y - 4 = 0
2 + y - 4 = 0
y - 2 = 0
y = 2
===========
Cheers,
Stan H.
=============